A circle centered at the origin whose radius is 2 units is inscribed in a right triangle whose base is 8 units and altitude 6 units. Find the equation of the circle.
Answer by Edwin McCravy(20054)


The triangle it's inscribed in has nothing to do with it.  
The equation of any circle is

(x-h)² + (y-k)² = r²

You are given the center as (h,k) = {0,0}, the origin, and you are given
the radius r = 2, so

(x-0)² + (y-0)² = 2²

or

x² + y² = 4

Why was the triangle even mentioned?  It is totally irrelevant.

Edwin